OpenStudy (anonymous):

Solve 6=1-2n+5 ahh help!

OpenStudy (texaschic101):

6 = 1 - 2n + 5 6 = 6 - 2n 6 - 6 = -2n 0 = -2n 0 = n check.. 6 = 1 - 2(0) + 5 6 = 6 - 0 6 = 6 (correct) n = 0